### v1

# A universal upper bound on the photon sphere radius in higher-dimensional black holes

Abstract
In this work, we derive a universal upper bound for the photon sphere radius in static, spherically symmetric, asymptotically flat black hole spacetimes of arbitrary dimension $n\geq 4$, in the presence of an anisotropic matter field satisfying the weak energy condition and a non-positive trace of the energy-momentum tensor. Using the effective potential method, we obtain the bound $r_{\gamma}\leq[(n-1)M]^{\frac{1}{n-3}}$, where $M$ is the ADM mass of the black hole. This bound reduces to $r_{\gamma}\leq 3M$ in four dimensions, consistent with the known result in the literature. Our result provides a dimension-dependent upper bound for photon spheres and deepens the understanding of spacetime structure in higher-dimensional gravitational theories.
I Introduction
Black holes, as fundamental predictions of general relativity and key astrophysical objects, continue to be a central focus of theoretical and observational research. The first direct image of a black hole shadow, captured by the Event Horizon Telescope (EHT) EventHorizonTelescope:2019dse ; EventHorizonTelescope:2022wkp , has spectacularly confirmed the existence of these compact objects and our understanding of strong-field gravity. Among the characteristic features of black hole spacetimes, the photon sphere—a hypersurface on which massless particles can orbit the black hole on unstable circular null geodesics—plays a crucial role. It determines the boundary of the black hole shadow Perlick:2010zh ; Falcke:1999pj , is intimately connected to the characteristic quasi-normal modes of black holes Cardoso:2008bp ; Berti:2009kk , and constrains the spatial extent of matter fields (hair) outside the horizon Nunez:1996xv ; Hod:2011aa .
For the four-dimensional Schwarzschild black hole, the photon sphere is located at $r_{\gamma}=3M$. For more general, static, spherically symmetric, and asymptotically flat black holes surrounded by matter fields (hairy black holes), it was proven by Hod that the photon sphere radius satisfies the universal bound $r_{\gamma}\leq 3M$, provided the matter obeys the weak energy condition and its energy-momentum tensor has a non-positive trace Hod:2013jhd . This bound is saturated by the vacuum (bald) Schwarzschild black hole. Subsequent studies have explored related upper and lower bounds for photon spheres of compact stars Hod:2020pim ; Peng:2020shc ; Liu:2024odv ; Hod:2023jmx ; Song:2025fgq .
Given the significant interest in higher-dimensional theories of gravity, such as string theory and brane-world models, a natural and important question arises: does a similar universal bound exist for the photon sphere in higher-dimensional black hole spacetimes? Understanding how spacetime dimensionality affects fundamental structures like the photon sphere is essential for probing the geometry of higher-dimensional gravity and identifying potential observational signatures of extra dimensions.
In this work, we extend Hod’s analysis to arbitrary dimensions $n\geq 4$. Within a model-independent framework for static, spherically symmetric, asymptotically flat black holes with anisotropic matter, we derive the characteristic equation for the photon sphere. Under the assumptions of the weak energy condition and a non-positive trace for the energy-momentum tensor, we prove the existence of a photon sphere and establish the universal upper bound:
$r_{\gamma}\leq[(n-1)M]^{\frac{1}{n-3}}\;.$ (1)
In four dimensions ($n=4$), this reduces to $r_{\gamma}\leq 3M$, confirming Hod’s result Hod:2013jhd . Our bound is dimension-dependent and suggests that the presence of matter fields tends to pull the photon sphere inward compared to the corresponding vacuum (Tangherlini) black hole. This work generalizes a key four-dimensional result to higher dimensions, deepening our understanding of black hole structure in extended theories of gravity.
The paper is organized as follows. In Sec. II, we describe the general higher-dimensional black hole spacetime and the corresponding Einstein field equations for an anisotropic fluid. Sec. III contains our main analysis: we derive the photon sphere condition, prove its existence, and establish the universal upper bound under appropriate energy conditions. We conclude with a summary and discussion of implications in Sec. IV. Throughout, we use natural units with $G=c=1$.
II DESCRIPTION OF THE SYSTEM IN THE HIGHER-DIMENSIONAL BLACK HOLE
We begin by recalling the general metric ansatz for a spacetime $(\mathcal{M},g)$ that possesses the symmetry of a codimension-2 maximally symmetric space, which reads Cao:2019vlu
$g=h_{AB}(y)dy^{A}dy^{B}+r^{2}(y)\gamma_{ij}(z)dz^{i}dz^{j}\,$, (2)
where $A=1,2$, and $i=1,\cdots\,,n-2$, and $\gamma_{ij}dz^{i}dz^{j}$ is the metric of the codimension-2 maximally symmetric space $(\mathcal{K},\gamma)$ with a sectional curvature $k=0\,,\pm 1$. The two dimensional part of $(\mathcal{M},g)$ with coordinates $\{y^{A}\}$ has a Lorentz signature and can be denoted by $(M,h)$.
The connection coefficients of the metric (2) are
[…]
$\Gamma^{A}{}_{ij}=-r(D^{A}r)\gamma_{ij},\quad\Gamma^{i}{}_{jA}=\frac{D_{A}r}{r}\delta^{i}{}_{j}\;$, (4)
where $^{(2)}\Gamma$ denotes the connection coefficients computed from the two-dimensional Lorentzian metric $(M,h)$, $D_{A}$ is the covariant derivative compatible with $h_{AB}$ on $(M,h)$, and $\hat{\Gamma}^{i}{}_{jk}$ are connection coefficients of the codimension-2 maximally symmetric space $(\mathcal{K},\gamma)$.
The components of the Riemann tensor are
$R_{ABCD}={}^{(2)}R_{ABCD}\;$, (5)
$R_{AiBj}=-r(D_{A}D_{B}r)\gamma_{ij}\;,$ (6)
$R_{ijkl}=r^{2}(1-D_{A}rD^{A}r)(\gamma_{ik}\gamma_{jl}-\gamma_{il}\gamma_{jk})\;,$ (7)
and the components of the Ricci tensor are
$R_{AB}={}^{(2)}R_{AB}-(n-2)\frac{D_{A}D_{B}r}{r}\;,$ (8)
$R_{ij}=[-r(D^{A}D_{A}r)+(n-3)(1-D_{A}rD^{A}r)]\gamma_{ij}\;,$ (9)
while the Ricci scalar is
$R={}^{(2)}R-2(n-2)\frac{D^{A}D_{A}r}{r}+\frac{(n-2)(n-3)}{r^{2}}(1-D_{A}rD^{A}r)\;.$ (10)
Consequently, the Einstein tensor can be expressed as
$G^{A}{}_{B}=-\frac{(n-2)}{r}D^{A}D_{B}r-\frac{1}{2}\bigg[\frac{(n-2)(n-3)}{r^{2}}(1-D_{C}rD^{C}r)-\frac{2(n-2)}{r}D_{C}D^{C}r\bigg]h^{A}{}_{B}\;$, (11)
$G^{i}{}_{j}=\bigg\{-\frac{1}{2}{}^{(2)}R+(n-3)\frac{D_{A}D^{A}r}{r}+\bigg[(n-3)-\frac{(n-2)(n-3)}{2}\bigg]\frac{1-D_{A}rD^{A}r}{r^{2}}\bigg\}r^{2}\delta^{i}{}_{j}\;$, (12)
$G^{A}{}_{i}=0\;.$ (13)
In this work, we focus on the static, spherically symmetric, and asymptotically flat higher-dimensional black hole systems. The $n$-dimensional Einstein-Hilbert action with matter fields is given by
$S=\int d^{n}x\sqrt{-g}\bigg(\frac{R}{16\pi}+\mathcal{L}_{M}\bigg)\;,$ (14)
where $g$ is the determinant of the metric tensor, and $\mathcal{L}_{M}$ denotes the Lagrangian density of matter.
The metric in Ref. Hod:2013jhd can be easily generalized to a $n$-dimensional black hole spacetime and written as Lee:2025xdm
$ds^{2}=-e^{-2\delta(r)}\mu(r)dt^{2}+\mu(r)^{-1}dr^{2}+r^{2}d\Omega^{2}_{n-2}\;,$ (15)
where the metric functions $\delta(r)$ and $\mu(r)$ depend only on the areal coordinate $r$, and
$d\Omega^{2}_{n-2}=d\theta_{1}^{2}+\sin^{2}\theta_{1}d\theta_{2}^{2}+\cdots+\sin^{2}\theta_{1}\cdots\sin^{2}\theta_{n-3}d\theta_{n-2}^{2}\;,$ (16)
represents the line element of the unit (n2)(n-2)-sphere.
The regularity of the event horizon at $r=r_{H}$ imposes the boundary conditions
$\mu(r_{H})=0\quad\mathrm{with}\quad\mu^{\prime}(r_{H})\geq 0\;.$ (17)
Asymptotic flatness requires that as $r\rightarrow\infty$,
$\mu(r\rightarrow\infty)\rightarrow 1\quad\mathrm{and}\quad\delta(r\rightarrow\infty)\rightarrow 0\;.$ (18)
Here, we do not assume $\delta(r)=0$, so our results also apply to hairy black-hole configurations Volkov:1998cc ; Volkov:2016ehx .
Taking $T^{t}{}_{t}=-\rho$$T^{r}{}_{r}=p_{r}$ and $T^{\theta}{}_{\theta}=T^{\phi}{}_{\phi}=p_{t}$ , where $\rho$, $p$ and $p_{T}$ are identified as the energy density, radial pressure, and tangential pressure, respectively. The static diagonal energy-momentum tensor of an anisotropic fluid can be expressed as
$T^{\mu}{}_{\nu}=\mathrm{diag}(-\rho(r),\,p_{r}(r),\,\underbrace{p_{t}(r),\,p_{t}(r),\,\ldots,\,p_{t}(r)}_{n-2\ \text{terms}})\;.$ (19)
From the Einstein field equations $G^{\mu}{}_{\nu}=8\pi T^{\mu}{}_{\nu}$ and using Eq. (11), we obtain
$\mu^{\prime}=\frac{(n-3)(1-\mu)}{r}-\frac{16\pi r\rho}{(n-2)}\;,$ (20)
$\delta^{\prime}=-\frac{8\pi r(\rho+p_{r})}{(n-2)\mu}\;.$ (21)
where the prime denotes differentiation with respect to $r$. Substituting Eq. (17) into Eq. (20) yields
$\rho(r_{H})\leq\frac{(n-3)(n-2)}{16\pi r_{H}^{2}}\;,\quad p_{r}(r_{H})=-\rho(r_{H})\;.$ (22)
The mass $m(r)$ enclosed within a sphere of radius $r$ can be written as
$m(r)=\frac{1}{2}r_{H}^{n-3}+\frac{8\pi}{n-2}\int_{r_{H}}^{r}x^{n-2}\rho(x)dx\;,$ (23)
where $\frac{1}{2}r_{H}^{n-3}$ is the horizon mass $m(r_{H})$. From Eqs. (20) and (23), the relation between $\mu$ and $m(r)$ in $n$-dimensional spacetime can be expressed as
$\mu(r)=1-\frac{2m(r)}{r^{n-3}}\;.$ (24)
The condition of finite mass configuration characteristics implies
$\lim\limits_{r\to\infty}r^{n-1}\rho(r)=0\;.$ (25)
Substituting Eqs. (20) and (21) into the energy-momentum tensor conservation equation
$T^{\mu}_{r;\mu}=0\;,$ (26)
yields
$p_{r}^{\prime}=\frac{2T+(n-1)\rho-(n+1)p_{r}}{2r}-(\frac{n-3}{2r}+\frac{8\pi rp_{r}}{n-2})\frac{\rho+p_{r}}{\mu}\;,$ (27)
where
$T=-\rho+p_{r}+(n-2)p_{t}\;$ (28)
is the trace of the energy momentum tensor $T^{\mu}{}_{\nu}$. Eq. (27) reduces to the four-dimensional case in Ref. Hod:2013jhd .
III Upper bound on the radius of the photon sphere of black holes
We now analyze the photon sphere and derive an upper bound for its radius. Due to the spherical symmetry of the system, we restrict our attention to particles moving in the equatorial plane, i.e., all polar angles satisfy
$\theta_{1}=\theta_{2}=\cdots=\theta_{n-3}=\frac{\pi}{2}\;,\quad\theta_{n-2}=\phi\;,$ (29)
where $\phi$ is the azimuthal coordinates. The Lagrangian for null geodesics in the spacetime (15) is
$2\mathcal{L}=-e^{-2\delta(r)}\mu(r)\dot{t}^{2}+\frac{\dot{r}^{2}}{\mu(r)}+r^{2}\dot{\phi}^{2}=0\;.$ (30)
where a dot denotes differentiation with respect to an affine parameter. Since the Lagrangian is independent of $t$ and $\phi$, there are two conserved quantities: the energy $E$ and the angular momentum $\mathcal{L}$. From the Lagrangian (30), one derives the generalized momenta
$p_{t}=-e^{-2\delta(r)}\mu(r)\dot{t}=-E\;,$ (31)
$p_{\phi}=r^{2}\dot{\phi}=L\;,$ (32)
$p_{r}=\mu(r)^{-1}\dot{r}\;.$ (33)
Substituting Eqs. (31) and (32) into Eq. (30) yields
$\dot{r}^{2}=\mu\bigg(\frac{E^{2}}{e^{-2\delta}\mu}-\frac{L^{2}}{r^{2}}\bigg)$ (34)
for the photon orbit. The corresponding effective potential can therefore be defined as
$V(r)=\mu\bigg(\frac{E^{2}}{e^{-2\delta}\mu}-\frac{L^{2}}{r^{2}}\bigg)\;.$ (35)
A photon sphere must satisfy the two conditions $V(r)=0$ and $V^{\prime}(r)=0$, which give
$-r\mu^{\prime}+2\mu(1+r\delta^{\prime})=0\;.$ (36)
Substituting Eqs. (20) and (21) into Eq. (36) leads to the characteristic equation
$\mathcal{R}(r_{\gamma})=0\;,$ (37)
for the photon sphere, where
$\mathcal{R}(r)=3-n+\mu(n-1)-\frac{16\pi r^{2}p_{r}}{n-2}\;.$ (38)
From Eq. (22), we obtain
$-p_{r}(r_{H})\leq\frac{(n-3)(n-2)}{16\pi r_{H}^{2}}\;.$ (39)
We first prove the existence of a photon sphere exterior to the horizon. At the horizon, using Eqs. (17), (38) and (39), we find
$\mathcal{R}(r_{H})\leq 0\;.$ (40)
Moreover, as $r\to\infty$, substituting Eqs. (18) into Eq. (38) together with the finite‑mass condition Eq. (25) gives
$\mathcal{R}(r\rightarrow\infty)=2>0\;.$ (41)
It is worth emphasizing that $\mathcal{R}(r)$ is a continuous function. Therefore, from Eqs. (40) and (41) we conclude that a photon sphere must exist in the region $r_{H}\leq r<\infty$. This implies that there is no photon sphere in the region $r_{H}\leq r<r_{\gamma}$ (where $r_{\gamma}$ denotes the innermost photon sphere), i.e.,
$\mathcal{R}(r_{H}\leq r<r_{\gamma})<0\;.$ (42)
To derive an upper bound, we follow the approach of Ref. Hod:2013jhd and introduce the pressure function $P(r)=r^{n}p_{r}(r)$. This construction, inspired by the four-dimensional case, will allow us to analyze the monotonicity of the radial pressure in a convenient form. Next, we shall deduce the behavior of $p_{r}$ by analyzing the properties of $P(r)$. Differentiating $P(r)$ with respect to $r$, we obtain
$P^{\prime}(r)=\frac{r^{n-1}}{2\mu}[\mathcal{R}(\rho+p_{r})+2\mu T]\;.$ (43)
It can be directly noted that this reduces to the four-dimensional case Hod:2013jhd .
We assume that the matter field outside the black hole event horizon satisfies the following conditions:
  • (1).
    Weak Energy Condition (WEC): The components of the energy-momentum tensor satisfy the weak energy condition (WEC). This means that the energy density of the matter fields is positive semidefinite, i.e.,
    $\rho\geq 0\;,$ (44)
    and that it bounds the pressures, which implies the inequality
    $\rho+p_{r}\geq 0\;,\quad\mathrm{and}\quad\rho+p_{t}\geq 0\;.$ (45)
  • (2).
    Non-positive Trace Condition: The trace of the energy-momentum tensor is assumed to be non-positive, this implies, i.e.,
    $T\leq 0\;.$ (46)
    This condition holds for many common fields, including electromagnetic fields and conformally invariant matter, and is a natural extension of the assumption used in the four-dimensional proof Hod:2013jhd .
From Eq. (22), one finds
$P(r_{H})=r^{n}p_{r}(r_{H})=-r^{n}\rho(r_{H})\;.$ (47)
Combining Eqs. (44) and (47), we find that the pressure function $P(r)$ satisfies
$P(r_{H})\leq 0\;.$ (48)
at the event horizon.
Next, we analyze the behavior of $P(r)$ in the region from the event horizon to the photon sphere. By substituting the photon sphere characteristic inequality (42), together with the energy conditions (45) and (28) for the matter field, into the pressure gradient (43), we obtain
$P^{\prime}(r_{H}\leq r<r_{\gamma})<0\;.$ (49)
This implies that the pressure function $P(r)$ is monotonically decreasing in the region $r_{H}\leq r<r_{\gamma}$. Then, combining Eqs. (48) and (49), we finds that $P(r)$ is a non-positive and monotonically decreasing function in the region $r_{H}\leq r<r_{\gamma}$, i.e.,
$P(r)<0\quad\mathrm{where}\quad r_{H}\leq r<r_{\gamma}\;.$ (50)
Accordingly, at the photon sphere, we have
$p(r_{\gamma})\leq 0\;.$ (51)
Substituting Eq. (51) into Eqs. (37) and (38), we get
$\mu(r_{\gamma})\leq\frac{n-3}{n-1}\;.$ (52)
From Eqs. (24) and (52), this yields the upper bound
$r_{\gamma}\leq[(n-1)m(r_{\gamma})]^{\frac{1}{n-3}}$ (53)
for the photon sphere. Since $m(r)$ is a non-decreasing function of $r$ and $m(r_{\gamma})\leq M$, where $M=m(r\to\infty)$ is the total ADM mass of the black hole spacetime, we finally obtain the universal upper bound
$r_{\gamma}\leq[(n-1)M]^{\frac{1}{n-3}}\;,$ (54)
In the four-dimensional case, this reduces to $r_{\gamma}\leq 3M$, which is consistent with the result in Ref. Hod:2013jhd . It is worth noting that this upper bound $3M$ is saturated by the photon sphere of the (bald) Schwarzschild black hole. Therefore, it is reasonable to hypothesize that the upper bound of the photon sphere radius is saturated by a bald black hole, and the presence of matter fields alters this radius, making it smaller than that of the hairless black hole.
It is worth emphasizing that, for static, spherically symmetric and asymptotically flat spacetimes of arbitrary dimension ($n\geq 4$) which satisfy the weak energy condition (45) and the non-positive trace condition (46), a universal upper bound for the photon sphere can be directly obtained from Eq. (54), without the need for further complicated calculations. Moreover, since the boundary of a black hole shadow is determined by the photon sphere, this result can also provide a theoretical benchmark for the observation of black hole shadows, potentially constraining extra dimensions.
IV Discussion and conclusion
In the context of general relativity and higher-dimensional gravitational theories, the photon sphere of a black hole, as a key structure influencing its shadow and dynamics, holds significant importance for establishing universal bounds on its radius. Although the relevant bounds in four-dimensional spacetime have been thoroughly discussed, it remains unclear whether similar universal upper bounds exist for the photon spheres of higher-dimensional black holes.
In this work, we perform an analysis of the photon spheres for static, spherically symmetric and asymptotically flat higher-dimensional black holes. In particular, under the constraints of the weak energy condition and the non-positive trace of the energy-momentum tensor, we obtain the upper bound $r_{\gamma}\leq[(n-1)M]^{\frac{1}{n-3}}$ for the radius of its photon sphere via an analytical method, where $n$ is the spacetime dimension and $M$ is the total ADM mass of the spacetime. It is worth noting that we have not restricted the specific form of $\mu(r)$ in the derivation process, therefore, this is a dimension-dependent universal upper bound. Moreover, the upper bound of the photon sphere radius for an arbitrary spacetime dimension can be directly obtained from Eq. (54), which simplifies the calculations involved in the conventional effective potential method.
Our result generalizes a fundamental geometric constraint from four to higher dimensions, strengthening the theoretical toolkit for analyzing black holes in theories beyond standard general relativity, such as string theory and brane-world models. The bound suggests that observational measurements of black hole shadow sizes, for instance by next-generation very-long-baseline interferometry, could in principle be used to test for signatures of extra dimensions by comparing the inferred photon sphere radius with the dimension-dependent upper limit.
Future research could extend this work in several promising directions. A natural and challenging extension would be to study rotating (Kerr-like) higher-dimensional black holes, investigating whether a similar bound exists and how it depends on the spin parameter. Exploring photon sphere bounds in asymptotically de Sitter or anti-de Sitter spacetimes would also be valuable, especially for applications in the gauge/gravity duality and cosmological contexts. Furthermore, understanding the precise saturation condition of the bound in higher dimensions and its relation to specific matter models or no-hair theorems presents an interesting theoretical challenge. Finally, further investigation into the relationship between the photon sphere bound and other black hole characteristics, such as quasinormal mode spectra, thermodynamic stability, or holographic complexity, could reveal deeper geometric principles.
In conclusion, we have derived a universal, dimension-dependent upper bound for the photon sphere radius in a broad class of higher-dimensional black holes. This work not only consolidates our understanding of black hole geometry across dimensions but also provides a new theoretical tool for probing the structure of spacetime in extended theories of gravity.